# VideoLLM Knows When to Speak: Enhancing Time-Sensitive Video Comprehension with Video-Text Duet Interaction Format

### v1

@@ 5.1 Dense Captioning
We use Shot2Story [7], a video-text dataset with segment-level captions, as our dense captioning training data. Specifically, we use the 43k human-annotated subset due to its high-quality and detailed annotations. We preprocess the data to serve our purposes, and an illustration of reformatting the video segment and caption annotations to video-text duet interaction format is in Fig. 2.
@@
We also create labels for the informative head in dense captioning tasks. According to the previous paragraph, the model can not have a comprehensive understanding of this video segment until it has viewed a sufficient portion of the segment (50% in this case). Meanwhile, once the caption has been generated as model response, we assume that the remaining frames in this video segment no longer provide new information that is not covered in the caption. Therefore, we set the informative head’s label to TRUE for frames between 50% of this segment and the insertion point of the response, and set labels to FALSE for the other frames, as shown in Fig 2.
To adapt to long video input, we also select videos with 2 to 4 minutes in length from COIN [27] as a dense captioning task to MMDuetIT. The annotations in COIN are reformatted using the same method as Shot2Story.
5.2 Multi-Answer Grounded Video QA
An important application scenario for the video-text duet interaction format is multi-answer grounded video question-answering (MAGQA). Consider when we are watching a live broadcast of a basketball game and want to track the actions of a particular player in the game. This exemplifies a MAGQA task: the question is ”What does this particular player do in the video?”. When this player performs an action, the model should respond with a description of this action (i.e., multiple answers) in a real time manner. We believe this newly proposed MAGQA task can be widely used in real-world scenarios when users interact with a live-streaming video.
We construct training data for this task using GPT4o-2024-08-06 [20]. Given the captions of all segments from the video as input, GPT4o is prompted to generate a question related to one or more captions. For each of the segment captions, if it is related to the question, then GPT4o should also generate an answer that can be inferred from this caption. Otherwise, GPT4o should reply with “Not Mentioned.”, and this answer is not added to the training data. We use the same insertion method of dense captioning task as described in Sec. 5.1, to insert the answers into the video stream and construct informative head labels, and the question is inserted at a random place before the first answer. We also use the same insertion method to convert the human-annotated Shot2Story test set and randomly sampled 2000 examples as the test set of our MAGQA benchmark in Sec. 6.3. Therefore, this dataset contains 36834 examples in the train set and 2000 examples in the test set. We name it as “Shot2Story-MAGQA-39k” and show its statistics in Tab. 1.
Data Quality Assessment
We sample 100 examples (with 290 answers) from our test set for manual quality assessment. Among the sampled examples, we find 1 example with a question unanswerable from the video, 5 examples have 6 answers (2.1%) that contradict the video content, and 5 examples have 7 answers (2.4%) unrelated to the question. Overall, manual quality assessment shows that above 95% data of our test set belongs to the high quality, which confirms the potential value of using Shot2Story-MAGQA-39k to benchmark models. The reason for the high quality is when the video captions are provided, generating questions and answers based on these text captions is a very simple task for advanced LLMs like GPT4o. However, we also find that in 21 examples, the video contains additional information that is not covered in the answers. This is because some questions are very general, like ”What scene is the video displaying?”, and describing scenes in videos elaborately has been a long-lasting challenge for annotating video datasets.
5.3 Temporal Video Grounding
We also add DiDeMo [8], $\text{HiREST}_{grounding}$ [38] and QuerYD [19], three temporal video grounding tasks in MMDuetIT. Note that these data are used only for training the relevance head, which is designed for performing temporal video grounding tasks and judging the relevance between the question and the video for QA tasks. The query is first added at the beginning of the input sequence. For frames that are annotated as relevant to the query, we set the relevance head’s label to TRUE; otherwise, we set it to FALSE.
5.4 Dataset Statistics
The data distribution of MMDuetIT is shown in Fig. 3. Note that this dataset only contains 109k examples, which is relatively small compared to modern post-training datasets like [17, 15, 31]. The reason is that due to computational resource constraints, we plan to demonstrate the feasibility of our proposed video-text duet interaction format by fine-tuning a state-of-the-art VideoLLM. We assume that the used backbone model already possesses enough video comprehension capabilities. By using a small dataset, we aim to train this model to efficiently adopt this new interaction with minimum catastrophic forgetting of its existing abilities.
num ex- amples answers per video words per ques./ans. video seg. len (sec)
Train 36834 2.96 7.75/12.17 4.22
Test 2000 3.04 7.77/12.17 4.28
Table 1: Dataset statistics of Shot2story-MAGQA-39k.
Figure 3: Data Distribution of MMDuetIT.
QVHighlights Charades-STA
mAP/HIT@1 R@0.5/0.7
TimeChat 14.5/23.9 32.2/13.4
VTimeLLM - 31.2/11.4
HawkEye - 31.4/14.5
VTG-LLM 16.5/33.5 33.8/15.7
LLaVA-OV-TC 17.6/32.9 33.1/12.4
LLaVA-OV-VT 19.0/40.0 36.5/12.3
MMDuet (Ours) 31.3/49.6 42.4/18.0
Table 2: Zero-shot performance on highlight detection and temporal video grounding.
YouCook2
SODAc/CIDEr/F1
TimeChat 1.2/3.4/12.6
VTG-LLM 1.5/5.0/17.5
LLaVA-OV-TC 1.9/3.3/21.8
LLaVA-OV-VT 2.5/6.7/14.0
MMDuet (Ours) 2.4/5.7/19.2
+ rm. prev. resp. 2.9/8.8/21.7
Table 3: Zero-shot performance on dense video captioning.
6 Experiments
Implementations
MMDuet is initialized with LLaVA-OneVision [15]. We train the model on MMDuetIT for one epoch. The training takes about one day on a node with 8 Tesla V100 GPUs, and the inference runs on 1 Tesla V100 GPU. Implementation details are listed in the appendix.
Baselines
As MMDuet mainly focuses on time-sensitive video tasks, we use the following baselines that are able to represent time spans in videos by different representation formats: TimeChat [23], VTimeLLM (7B) [10], HawkEye [31] and VTG-LLM [6]. Since the initialization of MMDuet is stronger than that of the baselines, for a fair comparison we also conduct a controlled experiment in which the only difference is the interaction format. Specifically, we use the same initialization model (LLaVA-OneVision), training data (MMDuetIT) and training schedule, but reformat the data to the respective interaction formats and video segment representation formats used by TimeChat and VTimeLLM to train two baseline models. We refer to these models as LLaVA-OV-TC and LLaVA-OV-VT.
6.1 Highlight Detection and Temporal Video Grounding
We use highlight detection and temporal video grounding to evaluate the performance of the relevance head of MMDuet. Baseline models are required to generate a list of float numbers to represent the relevance score for each clip in QVHighlights [14], and a start and end time for the relevant video span in Charades-STA. However, for LLaVA-OV-TC and LLaVA-OV-VT, despite using different prompts as input, we were still unable to instruct the model to output a sequence of scores as in [23]. Therefore, we follow the method of Charades-STA to instruct the model to output a related span, and assign the score to 1 for clips within this span and 0 otherwise. MMDuet uses the relevance score min-max normalized to $[0,1]$ as the score in QVHighlights, and to classify whether this frame is relevant and calculate frame-level IoU in Charades-STA.
Since the relevance head provides a relevance score immediately after each frame, its prediction cannot leverage the context from subsequent video frames. To mitigate this limitation, we smooth the relevance score sequence. Specifically, we set each frame’s smoothed relevance score as the mean value of its original score, the relevance scores of the preceding $w$ frames and the following $w$ frames, where $w$ is the window size. We set $w=2$ for QVHighlights and $w=6$ for Charades-STA. Results are shown in Tab. 2. We observe that, compared to the baselines, MMDuet exhibits a significantly greater improvement in performance on QVHighlights. This indicates that traditional VideoLLMs struggle with generating a long sequence of relevance scores using a text-based form or identifying multiple related video segments in its text-based responses, whereas MMDuet’s approach of directly assigning relevance scores to each frame circumvents this issue.
Parameter sensitivity of $w$
Though the $w$ is empirically set for the results in Tab. 2, in Fig. 4 we show that within a fairly large range of $w$, MMDuet consistently outperforms all baseline models. It confirms that MMDuet is robust to these tasks with respect to the window size.
6.2 Dense Video Captioning
We test dense video captioning performance on YouCook2 [41], a challenging task that requires models to output the caption, start point and end point for about 8 steps in a minutes-long cooking video. Baseline models output the start time, end time and caption for each step in the text-based form. For MMDuet, since this task requires the model to continuously identify important actions from the video and output periodically, we employ a heuristic method to determine whether a model response should be output after each frame (need_response function in Sec. 4.2). We sum up the informative score for each frame as the video plays. When the sum reaches a threshold $s$ (we set $s=2$), the model generates a response right after this frame as the caption for that step, and then we reset the sum to 00 to start a new round of sum.
However, MMDuet cannot directly predict when a step starts or ends just by this video-text duet interaction format, as the model is unable to determine whether a frame is the beginning of a step without observing enough subsequent content. To get the start and end time for each step as required by this task, we adopt a simple workaround: we use the time of the previous response and the current response as the start time and end time for a step. If two adjacent steps have the same caption, we merge them into one step.
It has been a long-lasting problem that LLMs tend to repeat previously-generated content [35], and we find that this problem is especially severe in dense video captioning. It indicates that VideoLLMs are probably generating captions relying on text shortcuts rather than the video content. We have attempted common solutions such as repetition penalty [12], which though is still sub-optimal. Since the responses from MMDuet are separated across multiple turns, we find that simply removing previously generated turns from the context (“rm. prev. resp.” for short) by not appending their attention keys and values to the KV Cache alleviates this issue, leading to a significant improvement in performance.
As shown in Tab. 3, MMDuet does not show significant improvements on F1 metric, likely due to the simple solution we use to derive the start and end time based on responses. Even so, the CIDEr and CODA_c metric (inaccurate predicted time spans can have negative effects on these metrics) of MMDuet is still higher than all baselines, indicating that MMDuet outperforms baselines in terms of text quality, possibly due to its facilitation to information retrieval discussed in Sec. 1.
Parameter sensitivity of $s$
We report the performance with threshold $s$ from 1 to 3 in Fig. 5. Results show that the performance is quite robust across a wide range of $s$, and we simply choose $s=2$ in Tab. 3. Note that we can use different $s$ to suit various downstream tasks especially in such zero-shot setting.
Figure 5: Performance on dense video captioning with different $s$.
Model Real- Time? In-Span Score LLaMA/GPT # turns (w/o. / w/. dedup) time per example
Baselines

### v2

@@ 5.1 Dense Captioning
We use Shot2Story Han2023Shot2Story20KAN, a video-text dataset with segment-level captions, as our dense captioning training data. Specifically, we use the 43k human-annotated subset due to its high-quality and detailed annotations. We preprocess the data to serve our purposes, and an illustration of reformatting the video segment and caption annotations to video-text duet interaction format is in Fig.˜2: we randomly sample a position from 50% to 75% time duration for the corresponding video segment, and insert the caption at that position as a model response. We also create labels for the informative head in dense captioning tasks by setting the informative head’s label to TRUE for frames between 50% of this segment and the insertion point of the response, and set labels to FALSE for the other frames. To adapt to long video input, we also select videos with 2 to 4 minutes in length from COIN Tang2019COIN as a dense captioning task to MMDuetIT. The annotations in COIN are reformatted using the same method as Shot2Story. For more details about this data reformat process please refer to Section˜B.1.
5.2 Multi-Answer Grounded Video QA
An important application scenario for the video-text duet interaction format is multi-answer grounded video question-answering (MAGQA). Consider when we are watching a live broadcast of a basketball game and want to track the actions of a particular player in the game. This exemplifies a MAGQA task: the question is "What does this particular player do in the video?". Each time this player performs an action, the model should respond with a description of this action (i.e., multiple answers) in a real time manner. We believe this newly proposed MAGQA task can be widely used in real-world scenarios when users interact with a live-streaming video.
We construct training data for this task using GPT4o-2024-08-06 openai2024gpt4o. Given the captions of all segments from the video as input, GPT4o is prompted to generate a question related to one or more captions. For each of the segment captions, if it is related to the question, then GPT4o should also generate an answer that can be inferred from this caption. Otherwise, GPT4o should reply with “Not Mentioned.”, and this answer is not added to the training data. We use the same insertion method of dense captioning task as described in Section˜5.1, to insert the answers into the video stream and construct informative head labels, and the question is inserted at a random place before the first answer. We also use the same insertion method to convert the human-annotated Shot2Story test set and randomly sampled 2000 examples as the test set of our MAGQA benchmark in Section˜6.3. Therefore, this dataset contains 36834 examples in the train set and 2000 examples in the test set, and we name it as “Shot2Story-MAGQA-39k”.
We have manually checked its data quality, and details of this process are stated in Appendix˜A.
5.3 Temporal Video Grounding
We also add DiDeMo Hendricks2017LocalizingMI, $\text{HiREST}_{grounding}$ Zala2023HierarchicalVR and QuerYD Oncescu2021QUERYDAV, three temporal video grounding tasks in MMDuetIT. Note that these data are used only for training the relevance head, which is designed for performing temporal video grounding tasks and judging the relevance between the question and the video for QA tasks. The query is first added at the beginning of the input sequence. For frames that are annotated as relevant to the query, we set the relevance head’s label to TRUE; otherwise, we set it to FALSE.
5.4 Dataset Statistics
The data distribution of MMDuetIT is shown in Fig.˜3. Note that this dataset only contains 109k examples, which is relatively small compared to modern post-training datasets like Li2023MVBenchAC; Li2024LLaVAOneVisionEV; Wang2024HawkEyeTV. The reason is that due to computational resource constraints, we plan to demonstrate the feasibility of our proposed video-text duet interaction format by fine-tuning a state-of-the-art VideoLLM. We assume that the used backbone model already possesses enough video comprehension capabilities. By using a small dataset, we aim to train this model to efficiently adopt this new interaction with minimum catastrophic forgetting of its existing abilities.
Figure 3: Data Distribution of MMDuetIT.
QVHighlights Charades-STA YouCook2
mAP/HIT@1 R@IoU=0.5/0.7 SODAc/CIDEr/F1
Video-LLaMA 11.3/15.6 2.7/1.2 0.0/0.0/0.1
VideoChat-Embed 13.1/18.1 3.2/1.4 0.2/0.6/3.4
VideoChatGPT - 7.7/1.7 -
TimeChat 14.5/23.9 32.2/13.4 1.2/3.4/12.6
VTimeLLM - 31.2/11.4 -
HawkEye - 31.4/14.5 -
VTG-LLM 16.5/33.5 33.8/15.7 1.5/5.0/17.5
VideoLLM-Online - - 0.4/0.9/5.8
LLaVA-OV-TC 17.6/32.9 33.1/12.4 1.9/3.3/21.8
LLaVA-OV-VT 19.0/40.0 36.5/12.3 2.5/6.7/14.0
MMDuet (Ours) 31.3/49.6 42.4/18.0 2.4/5.7/19.2
+ rm. prev. resp. - - 2.9/8.8/21.7
Table 1: Zero-shot performance on highlight detection, temporal video grounding, and dense video captioning. All models uses 7B LLMs.
6 Experiments
Implementations
MMDuet is initialized with LLaVA-OneVision Li2024LLaVAOneVisionEV. We train the model on MMDuetIT for one epoch. The training takes about one day on a node with 8 Tesla V100 GPUs, and the inference runs on 1 Tesla V100 GPU. More implementation details are listed in Section˜B.2.
Baselines
As MMDuet mainly focuses on time-sensitive video tasks, we use the following baselines that are able to represent time spans in videos by different representation formats: TimeChat Ren2023TimeChatAT, VTimeLLM (7B) Huang2023VTimeLLMEL, HawkEye Wang2024HawkEyeTV, VTG-LLM Guo2024VTGLLMIT, and VideoLLM-Online Chen2024VideoLLMonlineOV. For VideoLLM-Online, we experimented with $\theta\in\{0.5,0.6,0.7,0.8\}$ as suggested in their paper and report the best results (0.8 for both dense video captioning and MAGQA).
Since the initialization of MMDuet is stronger than that of the baselines, for a fair comparison we also conduct a controlled experiment in which the only difference is the interaction format. Specifically, we use the same initialization model (LLaVA-OneVision), training data (MMDuetIT) and training schedule, but reformat the data to the respective interaction formats and video segment representation formats used by TimeChat and VTimeLLM to train two baseline models. We name these models as LLaVA-OV-TC and LLaVA-OV-VT.
6.1 Highlight Detection and Temporal Video Grounding
We use highlight detection and temporal video grounding to evaluate the performance of the relevance head of MMDuet. Baseline models are required to generate a list of float numbers to represent the relevance score for each clip in QVHighlights Lei2021QVHighlightsDM, and a start and end time for the relevant video span in Charades-STA. However, for LLaVA-OV-TC and LLaVA-OV-VT, despite using different prompts as input, we were still unable to instruct the model to output a sequence of scores as in Ren2023TimeChatAT. Therefore, we follow the method of Charades-STA to instruct the model to output a related span, and assign the score to 1 for clips within this span and 0 otherwise. MMDuet uses the relevance score min-max normalized to $[0,1]$ as the score in QVHighlights, and to classify whether this frame is relevant and calculate frame-level IoU in Charades-STA.
Since the relevance head provides a relevance score immediately after each frame, its prediction cannot leverage the context from subsequent video frames. To mitigate this limitation, we smooth the relevance score sequence. Specifically, we set each frame’s smoothed relevance score as the mean value of its original score, the relevance scores of the preceding $w$ frames and the following $w$ frames, where $w$ is the window size. We set $w=2$ for QVHighlights and $w=6$ for Charades-STA. Results are shown in Table˜1. We observe that, compared to the baselines, MMDuet exhibits a significantly greater improvement in performance on QVHighlights. This indicates that traditional VideoLLMs struggle with generating a long sequence of relevance scores using a text-based form or identifying multiple related video segments in its text-based responses, whereas MMDuet’s approach of directly assigning relevance scores to each frame circumvents this issue. For VideoLLM-Online, we instruct it to reply with “start” / “end” at the start / end time of the target clip following the examples given in its paper but it does not follow the instructions despite trying different wordings, so we are not able to report its performance.
$w$ is robust to different values
Though the $w$ is empirically set for the results in Table˜1, we also find that within a fairly large range of $w$, the performance of MMDuet is robust and consistently outperforms all baseline models. Detailed results are listed in Section˜C.1.
6.2 Dense Video Captioning
We test dense video captioning performance on YouCook2 Zhou2017TowardsAL, a challenging task that requires models to output the caption, start point and end point for about 8 steps in a minutes-long cooking video. Baseline models output the start time, end time and caption for each step in the text-based form. For MMDuet, since this task requires the model to continuously identify important actions from the video and output periodically, we employ a heuristic method to determine whether a model response should be output after each frame (need_response function in Section˜4.2). We sum up the informative score for each frame as the video plays. When the sum reaches a threshold $s$ (we set $s=2$), the model generates a response right after this frame as the caption for that step, and then we reset the sum to 0 to start a new round of sum.
However, MMDuet cannot directly predict when a step starts or ends just by this video-text duet interaction format, as the model is unable to determine whether a frame is the beginning of a step without observing enough subsequent content. To get the start and end time for each step as required by this task, we adopt a simple workaround: we use the time of the previous response and the current response as the start time and end time for a step. If two adjacent steps have the same caption, we merge them into one step. This workaround is also applied on VideoLLM-Online.
It has been a long-lasting problem that LLMs tend to repeat previously-generated content Xu2022LearningTB, and we find that this problem is especially severe in dense video captioning. It indicates that VideoLLMs are probably generating captions relying on text shortcuts rather than the video content. We have attempted common solutions such as repetition penalty Keskar2019CTRLAC, which though is still sub-optimal. Since the responses from MMDuet are separated across multiple turns, we find that simply removing previously generated turns from the context (“rm. prev. resp.” for short) by not appending their attention keys and values to the KV Cache alleviates this issue, leading to a significant improvement in performance. However, this simple trick is not applicable to “whole-video” format baselines, as if the latest words are removed from the KV Cache, it will remain the same as before generating the latest words and the model will generate the same words again, despite some minor changes due to random sampling. In contrast, for MMDuet new video contents continuously bring new KV Cache and drive the conversation forward.
As shown in Table˜1, MMDuet does not show significant improvements on F1 metric, likely due to the simple solution we use to derive the start and end time based on responses. Even so, the CIDEr and CODA_c metric (inaccurate predicted time spans can have negative effects on these metrics) of MMDuet is still higher than all baselines, indicating that MMDuet outperforms baselines in terms of text quality, possibly due to its facilitation to information retrieval discussed in Section˜1.
$s$ is robust to different values
We also find that the threshold $s$ is quite robust across a wide range of from 1 to 3, and we can use different $s$ to suit various downstream tasks especially in such zero-shot setting. Detailed results are listed in Section˜C.1.
Model
Real-
Time?
original test set 5-time prolonged video test set
In-Span Score
LLaMA/GPT
# turns (w/o.
/ w/. dedup)
time per
example
In-Span Score
LLaMA/GPT
# turns (w/o.
/ w/. dedup)
time per
example
Baselines
LLaVA-OV-TC
2.77/2.64 4.1/2.2 1.00 1.67/1.62 7.6/2.4 1.00
LLaVA-OV-VT
2.54/2.42 4.1/3.1 1.06 1.64/1.60 10.2/3.4 0.99
VideoLLM-Online
1.33/1.26 1.3/1.1 0.44 - - -
MMDuet (Ours)
$t=0.6$
2.46/2.33 13.7/4.0 1.90 1.83/1.73 22.3/7.0 1.04
$t=0.5$
2.77/2.61 18.4/5.3 2.36 2.16/2.02 31.2/9.8 1.45
$t=0.4$
3.00/2.81 23.0/6.6 2.75 2.44/2.28 41.7/13.0 2.17
$t=0.3$
3.13/2.93 27.0/7.6 2.90 2.63/2.45 52.8/16.5 2.62
Table 2: Results on the test set of Shot2Story-MAGQA-39k with the rm. ass. turns method used. For the “time per example” column, the time used by “LLaVA-OV-VT” is set to 1, and the times for other rows are set as multiples of the time used by “LLaVA-OV-TC”. : Inference time of VideoLLM-Online is changed to gray and de-emphasized as it only generates one reply immediately after the question and is hardly helpful for answering the question, and thus we no longer evaluate it on the 5-times prolonged video test set.
6.3 Multi-Answer Grounded Video QA
To align closely with the widely-used streaming video comprehension scenario, we propose MAGQA that requires a model to generate answers at multiple necessary positions of a video. Different from conventional Video QA in which one question corresponds to only one answer, In MAGQA, a question corresponds to multiple turns of answers, and these turns are derived from different video segments. Therefore, this task requires the response to be accurate and in time. Though under the video-text duet interaction format users may raise arbitary number of questions at any time, to ensure the feasibility of evaluation, in this experiment we assume that the user raises only one question at the beginning of the video, and leave the extension to multiple questions as future work.
As this task is a newly-proposed one, we introduce an “in-span score” metric, which uses LLMs to calculate the average similarity of pred answers and gold answers that falls into the same time span of response, to evaluate both the correctness and timeliness of model responses. A detailed description of this metric is in Section˜B.5. To prevent reproducibility issues due to potential changes of OpenAI API, besides GPT-4o-2024-08-06 openai2024gpt4o, we also report the in-span score obtained using LLaMA 3.1 70B Instruct Dubey2024Llama3 to calculate pred-gold similarities.
As MAGQA requires the answers to be both informative and related to the question, we set need_response as: if the sum of informative score and relevance score of a frame is larger than a threshold $t$, then the model needs to generate a response right after this frame. We also use the “rm. prev. resp.” method in dense video captioning task introduced in Section˜6.2. As baseline models are not capable of generating responses at specific positions in the video, we employ an output format the same as dense video captioning, i.e., output the start time, end time, and predicted text for each turn after watching the entire video in both training and testing, and use the average of the start and end time as the response time. We also observe that for some cases the baseline models directly give one answer instead of generating multiple replies and their corresponding time spans, and we do not count these examples into the metrics when reporting results. Note that this is a significantly simplified requirement than that of MMDuet, as the MAGQA task simulates streaming video comprehension application scenario, which requires the model to respond as soon as the video plays to segments relevant to the question, which ensures that users can see the responses timely, rather than waiting until the entire video concludes before generating replies.
MMDuet has better performance than baselines and provides real-time replies.
Results on the test set of Shot2story-MAGQA-39k are shown in the lelf half of Table˜2. We provide results for different $t$ as it represents a trade-off between inference time and performance: as $t$ decreases from $0.6$ to $0.3$, the performance of MMDuet’s real-time replies continuously rises and outperforms baselines with a simplified setting of providing non-real-time replies after watching the entire video. However, this is achieved at a cost of generating lots of duplicate replies with more inference time.
MMDuet performs much better than baselines on longer videos.
Since the average video length of the test set of Shot2story-MAGQA-39k is only 16.9 seconds, to demonstrate MMDuet ’s real-time QA capabilities on longer videos we use a simple approach to make videos in the test set longer: we splice the video with 4 other videos randomly selected from the test set in random order to prolong the video to approximately 5 times longer by padding with videos irrelevant to the question. Results on the prolonged videos are shown in the right half of Table˜2. When the videos are long, it becomes harder for baseline models to output correct time spans for the answers which results in low in-span scores, while MMDuet is more likely to generate correct answers at the right time.
Model Acc
Flash-VStream 1.96
VLLM-Online 3.92
Dispider 25.34
MMDuet 29.44
Table 3: Performance on the Proactive Output task of StreamingBench.
Model YouCook2
MMDuet 2.9/8.8/21.7
w/o rand.
resp. pos.
2.1/7.3/19.0
w/o multi
informative
2.9/8.0/16.5
Table 4: Ablation study on training methods.
6.4 Proactive Output on StreamingBench
To further demonstrate the timeliness of the replies of MMDuet, we also report results on the Proactive Output task of StreamingBench Lin2024StreamingBenchAT. StreamingBench evaluates VideoLLMs in real-time, streaming video understanding tasks. Specifically, for the “Proactive Output” task, a question is considered as correctly answered if a reply is raised by the model within two seconds when a certain scene that contains the answer appears. Results in Table˜4 show that MMDuet outperforms all Streaming or Proactive MLLMs Zhang2024FlashVStreamMR; Chen2024VideoLLMonlineOV; Qian2025DispiderEV. Refer to Section˜C.2 for more details and baselines.
6.5 Ablation Studies
We conduct ablation studies on YouCook2 dense video captioning to assess two empirical yet important findings for effectively training the informative head in data construction: randomly inserting the response at a position from 50% to 75% of the corresponding video segment (rand. resp. pos.), and setting informative head’s label to TRUE for all frames between 50% of the segment and the response time (multi informative). When “rand. resp. pos.” is disabled, the response is always inserted at the end of the corresponding segment. When “multi informative” is disabled, only the informative label of the frame right before the response is set as TRUE. As illustrated in Table˜4, disabling either method negatively impact MMDuet’s performance, which shows the importance of carefully handling the response time and informative labels.
7 Conclusion
In this paper, we first formalize the video-text duet interaction format. We collect MMDuetIT for training models to follow the video-text duet interaction format. Based on MMDuetIT we train MMDuet, a model with significant improvements on various time-sensitive tasks and is able to automatically decide when to response in a real-time manner. We believe such improvements can be a substantial step towards building powerful and useful video comprehension applications.
Limitations
We acknowledge that there is much room for improvement which should be addressed in future research: (1) Some hyperparameters (e.g., the need_response criterion) are required during inference. However, we have shown that this criterion is quite robust across different thresholds. (2) Information from subsequent frames is not incorporated when generating in-time responses for the current frame, especially for the live-streaming video that indeed has unpredictable future frames. It can be crucial in some scenarios, such as determining the start of an action. (3) Slow inference speed. A better inference process is needed for avoid generating duplicate responses. (4) Real-time response datasets with longer live-streaming videos are required to be collected to better fit the real-world application scenarios.
Acknowledgement
This work is supported in part by the State Key Laboratory of General Artificial Intelligence.
Appendix A Data Quality Check of Shot2Story-MAGQA-39k
We sample 100 examples (with 290 answers) from our test set for manual quality assessment. Among the sampled examples, we find 1 example with a question unanswerable from the video, 5 examples have 6 answers (2.1%) that contradict the video content, and 5 examples have 7 answers (2.4%) unrelated to the question. Overall, manual quality assessment shows that above 95% data of our test set belongs to the high quality, which confirms the potential value of using Shot2Story-MAGQA-39k to benchmark models. The reason for the high quality is when the video captions are provided, generating questions and answers based on these text captions is a very simple task for advanced LLMs like GPT4o. However, we also find that in 21 examples, the video contains additional information that is not covered in the answers. This is because some questions are very general, like "What scene is the video displaying?", and describing scenes in videos elaborately has been a long-lasting challenge for annotating video datasets.
Appendix B Details of Training and Inference
B.1 Data Reformat Process of MMDuetIT
In Section˜5.1 we briefly introduced how the annotations for offline dense captioning / QA are converted into image-text interleave interactive format for training MMDuet. Here we elaborate more details and the reasons of this design:
@@ B.1 Data Reformat Process of MMDuetIT
We also create labels for the informative head in dense captioning tasks. According to the previous paragraph, the model can not have a comprehensive understanding of this video segment until it has viewed a sufficient portion of the segment (50% in this case). Meanwhile, once the caption has been generated as model response, we assume that the remaining frames in this video segment no longer provide new information that is not covered in the caption. Therefore, we set the informative head’s label to TRUE for frames between 50% of this segment and the insertion point of the response, and set labels to FALSE for the other frames.
B.2 Training Hyperparameters
LLaVA-OneVision uses SigLIP-Large Zhai2023SigmoidLF as the vision encoder, and converts an image with $384\times 384$ into $24\times 24=576$ tokens. In the official settings of LLaVA-OneVision Li2024LLaVAOneVisionEV, when encoding videos, the visual tokens corresponding to each frame are spatially downsampled to $12\times 12=144$ tokens using a pooling operation with a size of 2. However, this number of tokens is also too large when training and inference with long videos. To address this, we further modified the pooling size to 4, resulting in $7\times 7=49$ tokens per frame.